# Disorder-induced proximate quantum spin ice phase in $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$

### v1

@@ I Introduction
The discovery of classical spin ice (CSI) phases in rare-earth pyrochlore compounds marks a major success in the search for exotic magnetic phases with potential applications in quantum computing Broholm et al. (2020) and magnetic storage technologies Yao et al. (2013). The ground state of CSI is characterized by the “two-in, two-out” ice rules for semi-classical magnetic moments residing on corner-sharing tetrahedra and pointing along the local $\langle 111\rangle$ directions. In the first established CSI compounds, $\text{Ho}_{2}\text{Ti}_{2}\text{O}_{7}$ Fennell et al. (2009) and $\text{Dy}_{2}\text{Ti}_{2}\text{O}_{7}$ den Hertog and Gingras (2000), this ground state results from dominant Ising anisotropy and the self-screening dipolar interaction, which together generate ferromagnetic nearest-neighbor interactions. The CSI phase possesses a massive ground state degeneracy analogous to the degeneracy of proton order in water ice. It also exhibits dipolar-like spin-spin correlations that manifest as pinch-point features in magnetic neutron scattering patterns Fennell et al. (2009).
The success of CSI has inspired the theoretical proposal of quantum spin ice (QSI) Savary and Balents (2012), where anisotropic exchange interactions between spins enable quantum tunneling between different ice-rule-obeying configurations, thereby stabilizing a U(1) quantum spin liquid (QSL) phase with emergent electrodynamic properties. The QSI phase supports both gapless, photon-like magnetic excitations and gapped topological defects known as “magnetic monopoles” (or “visons,” as referred to in this work) Chen (2016, 2017). These monopoles represent topological defects of the emergent vector gauge potential in the U(1) QSL and have no classical analogue.
@@
Rare earth pyrochlores have featured prominently in the search for QSI candidates. Notable examples include $\text{Yb}_{2}\text{Ti}_{2}\text{O}_{7}$ Ross et al. (2011); Scheie et al. (2020), the Ce-based compounds $\text{Ce}_{2}\text{Zr}_{2}\text{O}_{7}$  Gaudet et al. (2019); Gao et al. (2019); Smith et al. (2022, 2025), $\text{Ce}_{2}\text{Sn}_{2}\text{O}_{7}$ Sibille et al. (2019); Porée et al. (2025), and $\text{Ce}_{2}\text{Hf}_{2}\text{O}_{7}$ Porée et al. (2023), and the Pr-based compounds $\text{Pr}_{2}\text{Zr}_{2}\text{O}_{7}$  Kimura et al. (2013); Wen et al. (2017); Petit et al. (2016a); Martin et al. (2017) and $\text{Pr}_{2}\text{Hf}_{2}\text{O}_{7}$  Sibille et al. (2016); Anand et al. (2016); Sibille et al. (2018). The QSI candidacy of $\text{Pr}_{2}\text{Zr}_{2}\text{O}_{7}$  was initially supported by the observation of narrow pinch points in the elastic magnetic scattering, and over 90% of the spectral weight in an inelastic continuum devoid of pinch-point features Kimura et al. (2013); Petit et al. (2016a). In contrast, the reported single crystals of its sister compound $\text{Pr}_{2}\text{Hf}_{2}\text{O}_{7}$  exhibit predominantly elastic magnetic scattering, accompanied by a gapped excitation near $\hbar\omega=0.2$ meV Sibille et al. (2016, 2018). Subsequent studies Wen et al. (2017); Martin et al. (2017) have revealed that disorder plays a prominent role in determining the ultra-low-temperature magnetic behavior of Pr-based pyrochlore magnets. In these systems, the non-Kramers ground state doublets of Pr3+ ions Kimura et al. (2013); Anand et al. (2016); Princep et al. (2013) can be split by any perturbation that breaks the local $D_{3d}$ symmetry. This effect can be modeled as a site-dependent transverse field $h_{i}$ acting on each Pr3+ ion, described by the Hamiltonian
@@
The longitudinal pseudospin operator $\sigma_{i}^{z}$ is related to the local magnetic moment through $\bm{m}_{i}=\mu_{\mathrm{eff}}\sigma_{i}^{z}\hat{\bm{z}}_{i}$, where $\mu_{\mathrm{eff}}$ denotes the effective moment of the ground-state doublet, and $\hat{\bm{z}}_{i}$ is the unit vector along the local $\langle 111\rangle$ direction (following the convention of Ref. Ross et al. (2011)). The transverse pseudospin components $\sigma_{i}^{x}$ and $\sigma_{i}^{y}$ represent quadrupolar degrees of freedom Petit et al. (2016a) and do not contribute to the magnetic dipole moment Savary and Balents (2017); Benton (2018). Nearest neighbor ($n=1$) antiferromagnetic interactions $J_{1}>0$ between pseudospins (equivalent to ferromagnetic interactions between magnetic moments) stabilize the “two-in, two-out” spin-ice state. In Eqn. 1, we have omitted the nearest-neighbor transverse interaction that is symmetrically allowed in Pr-based systems Petit et al. (2016a); Martin et al. (2017). Throughout this paper, we assume that a finite pairwise interaction between transverse pseudospin components may also be present, but focus primarily on the influence of $J_{1}$ and $h_{i}$.
@@
The interplay of spin-ice correlations and disorder-induced transverse fields can yield a rich space of QSL and confined phases Röchner et al. (2016); Savary and Balents (2017); Benton (2018); Pardini et al. (2019). A qualitative phase diagram, based on Refs. Savary and Balents (2017); Benton (2018); Pardini et al. (2019), is presented in Fig. 1, where $\bar{h}$ denotes the spatial average of the site-dependent transverse field $h_{i}$, while $\delta h$ characterizes the width of its spatial distribution. Starting from the clean limit, a small uniform transverse field $h_{i}$ enhances quantum fluctuations and can help stabilize the QSI phase Savary and Balents (2017) by generating a “random ring-exchange” term that facilitates quantum tunneling within the spin-ice manifold and enables the emergence of gapped topological excitations (visons). Two instabilities can drive the system out of the QSI phase, corresponding to the condensation of either spinons (spin flip defects) or visons. When $\bar{h}\gtrsim h_{c}\approx 0.6J_{1}$, spinons condense and the system transitions into a paramagnetic phase consisting of site-localized singlets Röchner et al. (2016); Benton (2018).In contrast, a finite $\delta h$ introduces spatial inhomogeneity in the ring-exchange amplitudes and enhances vison dispersion. Beyond a critical threshold $\delta h_{c}$, the system undergoes a confinement transition where the visons condense. This favours ordering with a (100) propagation vector Chen (2016); however, long-range (100) order is frustrated by structural disorder and the system instead enters a glassy phase with frozen magnetic moments. For QSI candidates, the critical disorder variance $\delta h_{c}$ at vanishing average field $\bar{h}\rightarrow 0$ is expected to remain finite due to quantum fluctuations arising from intrinsic transverse interactions. In systems with finite $\delta h$, an additional “Griffiths phase” Griffiths (1969); Savary and Balents (2017) may emerge, in which rare spatial regions locally satisfy $\bar{h}>h_{c}$ and enter the paramagnetic phase, while the bulk of the system either remains in a fluctuating QSI regime (“Griffiths QSI phase”) or transitions into a frozen state due to disorder-induced confinement (“Griffiths spin-frozen phase”).
@@
To realize a QSI state experimentally, it is essential to characterize the phases proximate to the QSI regime and to establish where candidate materials lie within the phase diagram. Previous studies Wen et al. (2017); Benton (2018) place $\text{Pr}_{2}\text{Zr}_{2}\text{O}_{7}$  deep within the paramagnetic phase. In contrast, $\text{Pr}_{2}\text{Hf}_{2}\text{O}_{7}$  samples—which exhibit significantly less inelastic scattering Sibille et al. (2018) and clear signatures of spin freezing Sibille et al. (2016)—have not been analyzed within the framework shown in Fig. 1. Among the Pr-based pyrochlores, $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$  has long been considered a promising QSI candidate: results on powder samples suggest that $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$  is a promising QSI candidate, since spins remain fluctuating down to at least 0.2 K, with no evidence of long-range magnetic ordering Zhou et al. (2008); Ortiz et al. (2024). However, large single crystals of $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$  suitable for inelastic neutron scattering have not previously been available.
@@
In this work, we present the first experimental study of flux-grown single-crystal $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$, combining comprehensive neutron scattering and thermodynamic measurements. Our key findings are as follows. First, the magnetic diffuse scattering observed within the spin-ice regime at $T\sim 0.3$ K cannot be accounted for by the classical nearest-neighbor spin-ice model; instead, either quantum fluctuations or further-neighbor interactions must be invoked. Second, magnetic susceptibility and inelastic neutron scattering measurements reveal two well-separated dynamical timescales, consistent with theoretical predictions for QSI states Tomasello et al. (2019) and with previous powder-sample results Ortiz et al. (2024), thereby lending strong support to the QSI scenario. Finally, below $T_{f}\approx 0.15$ K, our flux-grown single crystals enter a disorder-induced spin-frozen state with incipient $(100)$ magnetic order, positioned in close proximity to the QSI phase. This behavior contrasts sharply with earlier reports on floating-zone $\text{Pr}_{2}\text{Zr}_{2}\text{O}_{7}$ samples, which remain firmly in the paramagnetic regime Wen et al. (2017); Benton (2018).
@@ II Methods
Inelastic neutron scattering measurements were performed on the Cold Neutron Chopper Spectrometer (CNCS) at the Spallation Neutron Source (SNS), ORNL. Data were collected at three temperatures: $T=0.05$ K, $T=0.5$ K, and $T=10$ K. For each temperature, measurements were carried out using two incident neutron energies, $E_{i}=0.99$ meV and $E_{i}=3.32$ meV, both using the high-flux chopper configurations. These settings provided energy resolutions of approximately 0.02 meV and 0.1 meV at the elastic line, respectively. For both the WAND2 and CNCS experiments, the sample was rotated about the vertical $1\bar{1}0]$ axis over a total range of $180^{\circ}$, using step sizes of $0.1^{\circ}$ and $1^{\circ}$, respectively.
To assess the nature and extent of structural disorder in our sample, powder neutron-diffraction measurements were performed on the Nanoscale-Ordered Materials Diffractometer (NOMAD) at SNS, ORNL. Data were collected from crushed single-crystal specimens with a total mass of $\sim$0.5 g at $T=100$ K and $T=300$ K. Rietveld refinements of the diffraction patterns and pair-distribution function (PDF) analysis were carried out using the TOPAS  Coelho (2018) and PDFgui Farrow et al. (2007) software packages, respectively.
@@ III Results
Figure 3: Magnetic diffuse scattering of $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$  from experiment (a,e) and simulations (b–d, f–h). Panels (a) and (e) show experimental data at $T=0.3$ K and $T=1$ K, respectively, with high-temperature backgrounds ($T=10$ K and $T=20$ K) subtracted. Data are integrated along $K\bar{K}0]$ over $K\in[-0.1,0.1]$ rlu. Simulations include: (b,c,f,g) the quantum spin ice model from Ref. Benton et al. (2012), with (b,f) assuming the quantum limit and perturbative transverse interactions with uniform ring exchange $K^{\mathrm{eff}}=0.075(1)$ K; (c,g) the classical spin ice model with nearest-neighbor interaction $J_{1}$ only; and (d, h) best-fit results for dipolar spin ice (DSI) model with the calculated dipolar interaction $D_{\mathrm{dip}}\approx 0.012$ K (following the convention in Ref.Paddison (2023)), the exchange interaction $J_{1}^{\mathrm{ex}}=0.25(1)$ K, and $J_{2}^{\mathrm{ex}}=-0.018(1)$ K (see Eq. 1 and Appendix. D for the conventions).
@@ III.2 Magnetic diffuse scattering
The single-crystal magnetic diffuse scattering measured at $T=0.3$ K and $T=1$ K is shown in Fig. 3(a) and 3(e), respectively. At $T=1$ K, the $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$ sample already exhibits a characteristic “pinch-point” pattern, indicative of emerging spin-ice correlations. At $T=0.3$ K [Fig. 3(a)], a more asymmetric pattern develops: compared to the nearest-neighbor CSI simulation Benton et al. (2012) [Fig. 3(c)], we observe suppression of the pinch points, anisotropic intensity along the $[001]$ and $[111]$ directions, and broader scattering near the (222) position. Hence, our $T=0.3$ K data cannot be well described using the nearest-neighbor CSI model. To investigate the origin of the observed wavevector dependence, we performed fits using two distinct models. First, we examined the QSI model of Ref. Benton et al. (2012) characterized by a uniform ring-exchange interaction with fitted value $K^{\mathrm{eff}}=0.075(1)$ K (defined in Ref. Benton et al. (2012)).
@@
Second, a classical model was implemented Paddison (2023) incorporating fixed dipolar interactions $D_{\mathrm{dip}}\approx 0.012$ K (following the convention in Ref.Paddison (2023)) and fitted exchange couplings $J_{1}^{\mathrm{ex}}=0.25(1)$ K, and $J_{2}^{\mathrm{ex}}=-0.018(1)$ K (see Appendix D and Figs. A4 and A5 for details and conventions). Both the quantum model [Fig. 3(b,f)] and the classical model with next-nearest neighbor exchange interactions [Fig. 3(d,h)] provide an excellent description of our data.
@@
On the basis of equal-time magnetic correlations alone, it is therefore not possible to disentangle the role of quantum dynamics from that of further-neighbor interactions. To resolve this ambiguity, we turn to the dynamical properties. Theory predicts Tomasello et al. (2019) that a QSI phase in $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$ should exhibit two well-separated timescales: a slow component, $\tau_{\mathrm{slow}}$, detectable by a.c. susceptibility, and a fast component, $\tau_{\mathrm{fast}}$, observable by inelastic neutron scattering. We therefore collected single-crystal data using each of these probes.
@@
Figure 4: Bulk characterization of flux-grown single-crystal $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$. (a) Temperature dependence of the real part of the a.c. susceptibility $\chi^{\prime}$ measured along the $[111]$ direction at various frequencies and zero dc field. Data are normalized to the dc susceptibility between 1.8–4 K (black curve). Inset (i): Temperature dependence of the “slow” spin relaxation times Tomasello et al. (2019) of $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$ extracted from (a), fitted using a damped harmonic oscillator model Topping and Blundell (2019); Ortiz et al. (2024). (b) Imaginary part $\chi^{\prime\prime}$ of the a.c. susceptibility under the same conditions. (c) Arrhenius fit ($f=f_{0}\exp(-E_{a}/T_{f})$) of the $\chi^{\prime\prime}$ peak frequency, yielding an activation energy $E_{a}=2.7(7)$ K. Frequencies used in (a,b) are indicated in the color bar. (d) Heat capacity as a function of temperature after subtracting the lattice contribution estimated from non-magnetic La2Sn2O7, compared with two models for the nuclear Schottky anomaly: (1) the Kimura model Kimura et al. (2013), using $\mu_{\mathrm{hyp}}^{\mathrm{Pr}}=1.9~\mathrm{\upmu_{B}}$; and (2) the Gopal model Gopal (1966); Ortiz et al. (2024), assuming evenly spaced levels with $\epsilon_{r}=0.011$ meV. Inset (ii): Magnetic entropy obtained by subtracting the Gopal model from (d) and integrating $C_{m}/T$, recovering the full Pauling entropy near 10 K.

### v2

@@ I Introduction
The discovery of classical spin ice (CSI) phases in rare-earth pyrochlore compounds marks a major success in the search for exotic magnetic phases with potential applications in quantum computing [1] and magnetic storage technologies [2]. The ground state of CSI is characterized by the “two-in, two-out” ice rules for semi-classical magnetic moments residing on corner-sharing tetrahedra and pointing along the local $\langle 111\rangle$ directions. In the first established CSI compounds, $\text{Ho}_{2}\text{Ti}_{2}\text{O}_{7}$ [3] and $\text{Dy}_{2}\text{Ti}_{2}\text{O}_{7}$ [4], this ground state results from dominant Ising anisotropy and the self-screening dipolar interaction, which together generate ferromagnetic nearest-neighbor interactions. The CSI phase possesses a massive ground state degeneracy analogous to the degeneracy of proton order in water ice. It also exhibits dipolar-like spin-spin correlations that manifest as pinch-point features in magnetic neutron scattering patterns [3].
The success of CSI has inspired the theoretical proposal of quantum spin ice (QSI) [5], where anisotropic exchange interactions between spins enable quantum tunneling between different ice-rule-obeying configurations, thereby stabilizing a U(1) quantum spin liquid (QSL) phase with emergent electrodynamic properties. The QSI phase supports both gapless, photon-like magnetic excitations and gapped topological defects known as “magnetic monopoles” (or “visons,” as referred to in this work) [6, 7]. These monopoles represent topological defects of the emergent vector gauge potential in the U(1) QSL and have no classical analogue.
@@
Rare earth pyrochlores have featured prominently in the search for QSI candidates. Notable examples include $\text{Yb}_{2}\text{Ti}_{2}\text{O}_{7}$ [8, 9], the Ce-based compounds $\text{Ce}_{2}\text{Zr}_{2}\text{O}_{7}$  [10, 11, 12, 13, 14, 15], $\text{Ce}_{2}\text{Sn}_{2}\text{O}_{7}$ [16, 17, 18], and $\text{Ce}_{2}\text{Hf}_{2}\text{O}_{7}$ [19, 20], and the Pr-based compounds $\text{Pr}_{2}\text{Zr}_{2}\text{O}_{7}$  [21, 22, 23, 24] and $\text{Pr}_{2}\text{Hf}_{2}\text{O}_{7}$  [25, 26, 27]. The QSI candidacy of $\text{Pr}_{2}\text{Zr}_{2}\text{O}_{7}$  was initially supported by the observation of narrow pinch points in the elastic magnetic scattering, and over 90% of the spectral weight in an inelastic continuum devoid of pinch-point features [21, 23]. In contrast, the reported single crystals of its sister compound $\text{Pr}_{2}\text{Hf}_{2}\text{O}_{7}$  exhibit predominantly elastic magnetic scattering, accompanied by a gapped excitation near $\hbar\omega=0.2$ meV [25, 27]. Subsequent studies [22, 24] have revealed that disorder plays a prominent role in determining the ultra-low-temperature magnetic behavior of Pr-based pyrochlore magnets. In these systems, the non-Kramers ground state doublets of Pr3+ ions [21, 26, 28] can be split by any perturbation that breaks the local $D_{3d}$ symmetry. This effect can be modeled as a site-dependent transverse field $h_{i}$ acting on each Pr3+ ion, described by the Hamiltonian
@@
The longitudinal pseudospin operator $\sigma_{i}^{z}$ is related to the local magnetic moment through $\bm{m}_{i}=\mu_{\mathrm{eff}}\sigma_{i}^{z}\hat{\bm{z}}_{i}$, where $\mu_{\mathrm{eff}}$ denotes the effective moment of the ground-state doublet, and $\hat{\bm{z}}_{i}$ is the unit vector along the local $\langle 111\rangle$ direction (following the convention of Ref. [8]). The transverse pseudospin components $\sigma_{i}^{x}$ and $\sigma_{i}^{y}$ represent quadrupolar degrees of freedom [23] and do not contribute to the magnetic dipole moment [29, 30]. Nearest neighbor ($n=1$) antiferromagnetic interactions $J_{1}>0$ between pseudospins (equivalent to ferromagnetic interactions between magnetic moments) stabilize the “two-in, two-out” spin-ice state. In Eqn. 1, we have omitted the nearest-neighbor transverse interaction that is symmetrically allowed in Pr-based systems [23, 24]. Throughout this paper, we assume that a finite pairwise interaction between transverse pseudospin components may also be present, but focus primarily on the influence of $J_{1}$ and $h_{i}$.
@@
The interplay of spin-ice correlations and disorder-induced transverse fields can yield a rich space of QSL and confined phases [31, 29, 30, 32]. A qualitative phase diagram, based on Refs. [29, 30, 32], is presented in Fig. 1, where $\bar{h}$ denotes the spatial average of the site-dependent transverse field $h_{i}$, while $\delta h$ characterizes the width of its spatial distribution. Starting from the clean limit, a small uniform transverse field $h_{i}$ enhances quantum fluctuations and can help stabilize the QSI phase [29] by generating a “random ring-exchange” term that facilitates quantum tunneling within the spin-ice manifold and enables the emergence of gapped topological excitations (visons). Two instabilities can drive the system out of the QSI phase, corresponding to the condensation of either spinons (spin flip defects) or visons. When $\bar{h}\gtrsim h_{c}\approx 0.6J_{1}$, spinons condense and the system transitions into a paramagnetic phase consisting of site-localized singlets [31, 30].In contrast, a finite $\delta h$ introduces spatial inhomogeneity in the ring-exchange amplitudes and enhances vison dispersion. Beyond a critical threshold $\delta h_{c}$, the system undergoes a confinement transition where the visons condense [29, 30]. This favours ordering with a (100) propagation vector [6]; however, long-range (100) order is frustrated by structural disorder and the system instead enters a glassy phase with frozen magnetic moments. For QSI candidates, the critical disorder variance $\delta h_{c}$ at vanishing average field $\bar{h}\rightarrow 0$ is expected to remain finite due to quantum fluctuations arising from intrinsic transverse interactions. In systems with finite $\delta h$, an additional “Griffiths phase” [33, 29] may emerge, in which rare spatial regions locally satisfy $\bar{h}>h_{c}$ and enter the paramagnetic phase, while the bulk of the system either remains in a fluctuating QSI regime (“Griffiths QSI phase”) or transitions into a frozen state due to disorder-induced confinement (“Griffiths spin-frozen phase”).
@@
To realize a QSI state experimentally, it is essential to characterize the phases proximate to the QSI regime and to establish where candidate materials lie within the phase diagram. Previous studies [22, 30] place $\text{Pr}_{2}\text{Zr}_{2}\text{O}_{7}$  deep within the paramagnetic phase. In contrast, $\text{Pr}_{2}\text{Hf}_{2}\text{O}_{7}$  samples—which exhibit significantly less inelastic scattering [27] and clear signatures of spin freezing [25]—have not been analyzed within the framework shown in Fig. 1. Among the Pr-based pyrochlores, $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$  has long been considered a promising QSI candidate: results on powder samples suggest that $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$  is a promising QSI candidate, since spins remain fluctuating down to at least 0.2 K, with no evidence of long-range magnetic ordering [34, 35]. However, large single crystals of $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$  suitable for inelastic neutron scattering have not previously been available.
@@
In this work, we present the first experimental study of flux-grown single-crystal $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$, combining comprehensive neutron scattering and thermodynamic measurements. Our key findings are as follows. First, the magnetic diffuse scattering observed within the spin-ice regime at $T\sim 0.3$ K cannot be accounted for by the classical nearest-neighbor spin-ice model; instead, either quantum fluctuations or further-neighbor interactions must be invoked. Second, magnetic susceptibility and inelastic neutron scattering measurements reveal two well-separated dynamical timescales, consistent with theoretical predictions for QSI states [36] and with previous powder-sample results [35], thereby lending strong support to the QSI scenario. Finally, below $T_{f}\approx 0.15$ K, our flux-grown single crystals enter a disorder-induced spin-frozen state with incipient $(100)$ magnetic order, positioned in close proximity to the QSI phase. This behavior contrasts sharply with earlier reports on floating-zone $\text{Pr}_{2}\text{Zr}_{2}\text{O}_{7}$ samples, which remain firmly in the paramagnetic regime [22, 30].
@@ II Methods
Inelastic neutron scattering measurements were performed on the Cold Neutron Chopper Spectrometer (CNCS) at the Spallation Neutron Source (SNS), ORNL. Data were collected at three temperatures: $T=0.05$ K, $T=0.5$ K, and $T=10$ K. For each temperature, measurements were carried out using two incident neutron energies, $E_{i}=0.99$ meV and $E_{i}=3.32$ meV, both using the high-flux chopper configurations. These settings provided energy resolutions of approximately 0.02 meV and 0.1 meV at the elastic line, respectively. For both the WAND2 and CNCS experiments, the sample was rotated about the vertical $1\bar{1}0]$ axis over a total range of $180^{\circ}$, using step sizes of $0.1^{\circ}$ and $1^{\circ}$, respectively. The data were reduced and processed using the Mantid software suite [37].
To assess the nature and extent of structural disorder in our sample, powder neutron-diffraction measurements were performed on the Nanoscale-Ordered Materials Diffractometer (NOMAD) at SNS, ORNL. Data were collected from crushed single-crystal specimens with a total mass of $\sim$0.5 g at $T=100$ K and $T=300$ K. Rietveld refinements of the diffraction patterns and pair-distribution function (PDF) analysis were carried out using the TOPAS  [38] and PDFgui [39] software packages, respectively.
@@ III Results
Figure 3: Magnetic diffuse scattering of $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$  from experiment (a,e) and simulations (b–d, f–h). Panels (a) and (e) show experimental data at $T=0.3$ K and $T=1$ K, respectively, with high-temperature backgrounds ($T=10$ K and $T=20$ K) subtracted. Data are integrated along $K\bar{K}0]$ over $K\in[-0.1,0.1]$ rlu. Simulations include: (b,c,f,g) the quantum spin ice model from Ref. [40], with (b,f) assuming the quantum limit and perturbative transverse interactions with uniform ring exchange $K^{\mathrm{eff}}=0.075(1)$ K; (c,g) the classical spin ice model with nearest-neighbor interaction $J_{1}$ only; and (d, h) best-fit results for dipolar spin ice (DSI) model with the calculated dipolar interaction $D_{\mathrm{dip}}\approx 0.012$ K (following the convention in Ref.[41]), the exchange interaction $J_{1}^{\mathrm{ex}}=0.25(1)$ K, and $J_{2}^{\mathrm{ex}}=-0.018(1)$ K (see Eq. 1 and Appendix. D for the conventions).
@@ III.2 Magnetic diffuse scattering
The single-crystal magnetic diffuse scattering measured at $T=0.3$ K and $T=1$ K is shown in Fig. 3(a) and 3(e), respectively. At $T=1$ K, the $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$ sample already exhibits a characteristic “pinch-point” pattern, indicative of emerging spin-ice correlations. At $T=0.3$ K [Fig. 3(a)], a more asymmetric pattern develops: compared to the nearest-neighbor CSI simulation [40] [Fig. 3(c)], we observe suppression of the pinch points, anisotropic intensity along the $[001]$ and $[111]$ directions, and broader scattering near the (222) position. Hence, our $T=0.3$ K data cannot be well described using the nearest-neighbor CSI model. To investigate the origin of the observed wavevector dependence, we performed fits using two distinct models. First, we examined the QSI model of Ref. [40] characterized by a uniform ring-exchange interaction with fitted value $K^{\mathrm{eff}}=0.075(1)$ K (defined in Ref. [40]).
@@
Second, a classical model was implemented [41] incorporating fixed dipolar interactions $D_{\mathrm{dip}}\approx 0.012$ K (following the convention in Ref.[41]) and fitted exchange couplings $J_{1}^{\mathrm{ex}}=0.25(1)$ K, and $J_{2}^{\mathrm{ex}}=-0.018(1)$ K (see Appendix D and Figs. A4 and A5 for details and conventions). Both the quantum model [Fig. 3(b,f)] and the classical model with next-nearest neighbor exchange interactions [Fig. 3(d,h)] provide an excellent description of our data.
@@
On the basis of equal-time magnetic correlations alone, it is therefore not possible to disentangle the role of quantum dynamics from that of further-neighbor interactions. To resolve this ambiguity, we turn to the dynamical properties. Theory predicts [36] that a QSI phase in $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$ should exhibit two well-separated timescales: a slow component, $\tau_{\mathrm{slow}}$, detectable by a.c. susceptibility, and a fast component, $\tau_{\mathrm{fast}}$, observable by inelastic neutron scattering. We therefore collected single-crystal data using each of these probes.
@@
Figure 4: Bulk characterization of flux-grown single-crystal $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$. (a) Temperature dependence of the real part of the a.c. susceptibility $\chi^{\prime}$ measured along the $[111]$ direction at various frequencies and zero dc field. Data are normalized to the dc susceptibility between 1.8–4 K (black curve). Inset (i): Temperature dependence of the “slow” spin relaxation times [36] of $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$ extracted from (a), fitted using a damped harmonic oscillator model [42, 35]. (b) Imaginary part $\chi^{\prime\prime}$ of the a.c. susceptibility under the same conditions. (c) Arrhenius fit ($f=f_{0}\exp(-E_{a}/T_{f})$) of the $\chi^{\prime\prime}$ peak frequency, yielding an activation energy $E_{a}=2.7(7)$ K. Frequencies used in (a,b) are indicated in the color bar. (d) Heat capacity as a function of temperature after subtracting the lattice contribution estimated from non-magnetic La2Sn2O7, compared with two models for the nuclear Schottky anomaly: (1) the Kimura model [21], using $\mu_{\mathrm{hyp}}^{\mathrm{Pr}}=1.9~\mathrm{\upmu_{B}}$; and (2) the Gopal model [43, 35], assuming evenly spaced levels with $\epsilon_{r}=0.011$ meV. Inset (ii): Magnetic entropy obtained by subtracting the Gopal model from (d) and integrating $C_{m}/T$, recovering the full Pauling entropy near 10 K.